Grade 11Algebra

in a sequence of 25 terms, the first 13 terms are in AP with the common difference 2 and the last 13 terms are in GP with the common ratio 2. if the middle term of the AP is equal to the middle term of the GP . THEN the middle term of the entire sequence is

KAUSHIK SUBRAHMANIAN
10 Years agoGrade 11

1 Answer

ronit
10 Years ago
it means that 
the middle term of the entire term=last term of the AP=first term of the GP
a1+12(2)=a2=>a1=a2-24
middle term of AP=middle term of GP
a1+6(2)=64a2=>a2-24+12=64a2=>63a2=-12=>a2=-12/63=>a2=-4/21
middle term of the entire sequence=-4/21
you can check it with the last term of AP
a1=(-4/21)-24=-508/21
last term of AP=(-508/21)+12(2)=(-508/21)+24=-4/21
still, i have a bit of doubt due to the negative answer with fraction